The ratio of two sides of two cubes is 2:3 and the difference of their volume is 152 cm^3. What is the length of the side of the bigger cube?
Answer by Alan3354(69443)

The ratio of two sides of two cubes is 2:3 and the difference of their volume is 152 cm^3. What is the length of the side of the bigger cube?
---------------
Vol = k*s^3 where k = constant and s = side length
--> ratio of volumes is 8:27
-----
27k - 8k = 152
k = 8
------
Volumes are 64 & 216
Sides are 4 & 6